OpenStudy (anonymous):

When comparing fifteen thirty-seconds to one-half using mental math, what fraction is useful? sixteen thirty-seconds fifteen-halves fifteen-thirtieths thirty-two sixty-fourths

OpenStudy (hoblos):

fifteen thirty-seconds = 15/32 one-half = 1/2 in order to compare you must have same denominator 1/2 = 16/32 and since 15/32 < 16/32 then 15/32 < 1/2 so you use sixteen thirty-seconds